AM4 -Esercitazione 10 A.A.2003-2004

Prof. Luigi Chierchia, Dott. Laura Di Gregorio 4 dicembre 2003

Esercizio

Si calcolino i coefficienti di Fourier delle seguenti funzioni periodiche di periodo 2π:

1. f(x) = e^{x} dispari ( 0 < x ≤ π);

2. f(x) = e^{x} su [−π, π);

3. f(x) = \frac{π − x}{2} su [0, 2π);

4. f(x) = ∑_{k=1}^{∞} \frac{sin^{2} k^{2}x}{2^{k}} su [−π, π).